# Fluid Mechanics Unit Italian Aerospace Research Center (CIRA) Capua (CE), Italy

### v1

@@ 4.1 On the kinetic-energy and pressure-equilibrium preservation of the new schemes
[…]
We note that the convective term in the kinetic energy equation is in a quadratically split form which will result in local conservation of the quantities. Analogous expressions can be written for the biased stencils. Although kinetic energy preservation does not provide a rigorous nonlinear stability metric for compressible flows, incorporating its numerical consistency has been shown to greatly enhance solution quality—even for the class of entropy conserving Tadmor fluxes [51, 27]. And as motivated in Eqs. (5) and (2.1) of Section 2, it forms a key consistency within the continuous entropy analysis.
Next, it is possible to also show that all the new FD-EC schemes have the Pressure Equilibrium Preservation (PEP) property [4, 26, 53] for the current context of a calorically-perfect single fluid. In other words, they have the ability to preserve the equilibrium of velocity and pressure across an interface. The condition on the discretization of the momentum equation assuming constant velocity $U_{o}$ and pressure $P_{o}$ (see [53, 26]) is that the momentum flux is equal to
@@ 5 Numerical results
The new EC schemes are here tested with a variety of cases (i.e., the 2D isentropic vortex, the 1D density wave, the 1D Sod shock tube, and the 3D Taylor Green vortex) in order to verify the validity of the theoretical results. In particular, we consider the new scheme based on the wide stencil presented in Eq. (77), here referred to as ECW, and three schemes from the family identified in Eq. (84), that with $\alpha=0$, $\alpha=1$, and $\alpha=0.5$ are labeled ECB, ECF, and ECS, respectively. Notably, ECSis analogous to the scheme of Ranocha in the scenario of a two-point second order stencil. A summary of the schemes and their respective numerical fluxes is presented in Table 3.
@@
Despite having demonstrated the capability to obtain high-order formulations of the presented EC schemes, the methods will all be tested with their low-order version, employing first and second order accurate differential operators. This is due to the singularity problem that occurs when evaluating the logarithmic mean for almost identical values. This challenge is handled in the case of two-point means by employing suitable expansion approximations such as the one famously introduced by Ismail and Roe [47]. Other similar approaches have been recently employed to produce schemes that are asymptotically entropy conserving [24, 28]. However, a generalization of the procedure for multiple-point means has yet to be developed.
[…]
Ref. $\mathcal{F}_{\rho}$ $\mathcal{F}_{\rho u_{m},i+1/2}$ $\mathcal{F}_{\rho E}$ ECB new $\overline{(\rho_{i},\rho_{i+1})}^{(\log)}\,u_{n,i}$ $\mathcal{F}_{\rho}\overline{u_{m}}\,+p_{i+1}$ $\mathcal{F}_{\rho}\overline{(e_{i},e_{i+1})}^{(H\log)}\,+\mathcal{F}_{\rho}\,\sum_{m}\frac{u_{m,i}u_{m,i+1}}{2}+p_{i+1}u_{n,i}$ ECF new $\overline{(\rho_{i},\rho_{i+1})}^{(\log)}\,u_{n,i+1}$ $\mathcal{F}_{\rho}\overline{u_{m}}\,+p_{i}$ $\mathcal{F}_{\rho}\overline{(e_{i},e_{i+1})}^{(H\log)}\,+\mathcal{F}_{\rho}\,\sum_{m}\frac{u_{m,i}u_{m,i+1}}{2}+p_{i}u_{n,i+1}$ ECS [51] $\overline{(\rho_{i},\rho_{i+1})}^{(\log)}\,\overline{u_{n}}\,$ $\mathcal{F}_{\rho}\overline{u_{m}}\,+\overline{p}\,$ $\mathcal{F}_{\rho}\overline{(e_{i},e_{i+1})}^{(H\log)}\,+\mathcal{F}_{\rho}\,\sum_{m}\frac{u_{m,i}u_{m,i+1}}{2}+\overline{\overline{(p,u)}}$ ECW new $0.5\overline{(\rho_{i-1},\rho_{i+1})}^{(\log)}\,u_{n,i}$ $\mathcal{F}_{\rho}\overline{u_{m}}\,+\overline{p}\,$ $0.5\overline{(\rho_{i-1},\rho_{i+1})}^{(\log)}\,u_{n,i}\overline{(e_{i-1},e_{i+1})}^{(H\log)}\,$ $+0.5\overline{(\rho_{i},\rho_{i+2})}^{(\log)}\,u_{n,i+1}$ $+0.5\overline{(\rho_{i},\rho_{i+2})}^{(\log)}\,u_{n,i+1}\overline{(e_{i},e_{i+2})}^{(H\log)}\,$ $+\mathcal{F}_{\rho}\,\sum_{m}\frac{u_{m,i}u_{m,i+1}}{2}+\overline{\overline{(p,u)}}$
@@
Table 3: Numerical fluxes for the schemes used for the tests. The fluxes are evaluated along direction $n$ at location $i+1/2$. All interpolation operation are implicitly interpreted to be in the same direction $n$.
@@
The isentropic vortex flow is an exact solution of the inviscid compressible flow equations and as such it can be used to evaluate the accuracy of the numerical methods. In Fig. 1(a) the error on momentum along the $x$ axis is displayed as a function of $N$ number of nodes in each direction. The error is evaluated as the maximum norm of the difference between the computed and the exact solution for momentum along the $x$ direction $\|\rho u-(\rho u)_{ex}\|_{\infty}$, evaluated at a time $t=0.01$ with $\text{CFL}=0.01$. The analysis shows a difference between the first-order accurate (ECFand ECB) and the second-order accurate (ECSand ECW) methods. The reason behind this behaviour is easily explained by the fact that ECFand ECBuse an asymmetrical numerical derivative for the discretization of the pressure term and an asymmetrical interpolator for velocity as well; on the other hand both ECSand ECWresult in central schemes. Identical trends are witnessed upon inspecting the error of other quantities (not shown), such as the primitive variables $(\rho,u,T)$.
@@
Figure 2: Density wave test for different discretizations: black lines are used for ECB, red for ECS, green for ECF, blue ECW; dashed magenta line is the exact solution. $\text{CFL}=0.01$, final solutions at $T=40$.
It can be shown theoretically (see Section 4.1) that all of the current schemes have the PEP property. The simulation of a traveling density wave is well suited to numerically test the PEP property of the schemes. The initial conditions for the test are
@@
with the domain $[-L,L]$ with $L=1$ discretized in $61$ points and periodic boundary conditions. The Courant number has been chosen to be $\text{CFL}=0.001$ to prevent errors on entropy conservation due to the temporal integrator, and the simulation has been carried out until a final time $T=40$, which corresponds to approximately 20 cycles of the wave through the domain.
@@
The simulation is carried out until $T=0.1$ with $\text{CFL}=0.01$ based on velocity $u_{\infty}=\sqrt{p_{U}/\rho_{U}}$. For the discretization of the interval $[-0.25,0.25]$ $101$ nodes are employed, however a much larger domain has been used to prevent the periodic boundary conditions to have any effect on the solution.
In Fig. 3, the solutions obtained with the various EC schemes are compared with the analytical one obtained with an exact Riemann solver. As expected, all the schemes are entropy conserving and generate spurious oscillations in the solution. As visible in Fig. 3(c), none of the schemes are able to reproduce the correct behavior for entropy in the region leading up to shock $x\in[0.1,0.2]$, and ECWseems to present more oscillations. This behavior has been observed before for the scheme of Ranocha [26] (here ECS) in [31]. This region is also one in which the oscillations are stronger for all the schemes when other variables are considered (e.g., density, velocity, temperature, pressure).
On the other hand, in the region $x\in[0,0.1]$ the asymmetric schemes ECBand ECFpresent oscillations with smaller amplitude when compared to the central schemes ECWand ECS. This is particularly evident when inspecting velocity in Fig. 3(b), for which a larger spike is present near $x=0$ for the ECSscheme. Such behavior further motivates interest in studying these biased schemes for use in flows exhibiting sharp gradients.
@@
(d) Temperature $T$
(e) Pressure $p$
(f) Entropy integral evolution
5.4 3D Taylor Green vortex
(a) Time evolution of entropy integral
(b) Time evolution of kinetic energy integral
Figure 4: Taylor-Green vortex test for different discretizations: black lines are used for ECB, red for ECS, green for ECF, blue ECW. On the left-hand side, global quantities are evaluated at each time step of the simulation; on the right-hand side, they are sampled every $200$ steps. $\text{CFL}=0.01$.
(a) Time evolution of density fluctuations
(b) Time evolution of temperature fluctuations
Figure 5: Taylor-Green vortex test for different discretizations: black lines are used for ECB, red for ECS, green for ECF, blue ECW. On the left-hand side, fluctuations are evaluated at each time step of the simulation; on the right-hand side, they are sampled every $200$ steps. $\text{CFL}=0.01$.
The inviscid Taylor Green vortex is used to test the schemes in a three-dimensional setting that serves as a benchmark for turbulence simulations; it generates smaller and smaller scales through vortex stretching. The initial conditions are given by
$\rho(x,y,z)$ $=1$
$u(x,y,z)$ $=\sin(x)\cos(y)\cos(z)$
$v(x,y,z)$ $=-\cos(x)\sin(y)\cos(z)$
$w(x,y,z)$ $=0$
$p(x,y,z)$ $=100+\frac{(\cos(2x)+\cos(2y))(\cos(2x)+2)-2}{16}$
which correspond to a Mach number $M\approx 0.08$. The domain is a triperiodic cube of length $2\pi$ and it is discretized using $32\times 32\times 32$ nodes. The final time of the simulation is $T=70$ and the time step size adapts to keep $\text{CFL}=0.01$.
Entropy is conserved for the test case (Fig. 4(a)), but, while for the central schemes ECWand ECSthe error is down to $10^{-14}$, for the asymmetric schemes ECBand ECFthere is a drift in the integral value. This accumulation of errors is not observed in the conservation of the primary variables density, momentum and total energy (not shown here). On the other hand, the global value of kinetic energy presents clear oscillations for ECBand ECFschemes, which is likely connected to the first order discretization of the pressure term that governs the exchange between kinetic and internal energy. Fig. 4(b) shows the evolution of global kinetic energy considering the value at each time step of the simulation (on the left) and, for additional clarity, sampling every $200$ steps (on the right) [^7] The under-sampled plots suggest a leading behavior for the ECFscheme after awhile; however, this is an artifact of the chosen sampling rate..
A similar behavior is also displayed for the evolution of density and temperature fluctuation $\rho^{\prime}$ and $T^{\prime}$ in Fig. 5. However, despite the oscillations, none of the schemes yield an unbounded growth of the thermodynamic fluctuations. This is a desirable outcome since, after an initial transient, the flow should behave like inviscid isotropic homogeneous turbulence [20, 21, 56]. This successful result should not to be taken for granted, as many discretizations fail this test and present uncontrolled production in the fluctuations [3, 57].
6 Conclusions
The current work introduces a family of entropy conserving schemes for a calorically-perfect single-component gas. The methods have the additional properties of also being kinetic-energy-preserving and pressure-equilibrium-preserving. These schemes are, to the best of our knowledge, the first finite-difference discretizations to be locally conservative in both the primary quantities (mass, momentum, total energy) as well as the discrete entropy dynamics. Their ability to attain exact discrete conservation is confirmed with a variety of test cases in one-, two- and three-dimensions; meanwhile, the propagation of a density wave is also simulated to test the PEP property of the schemes.
The new framework introduces and leverages FD representations of the logarithmic mean, which is commonly employed in EC numerical fluxes. This equivalence is just one instance of a more general link between linear and nonlinear two-point interpolants and FD forms involving the ratio of differentials. Some of the most commonly used means (e.g., arithmetic, geometric, harmonic) are also provided in this paper. While the schemes presented in this work are initially introduced in the context of finite differencing, they are also locally conservative due to their SBP property and therefore admit a local flux form that can be used in a FV or FE setting. Even though we considered uniform grids herein, extension to arbitrary nodal distributions is also possible as long as a diagonal-norm SBP property is available. The corresponding EC property of the methods is based on point-wise cancellations of spurious volumetric terms that would otherwise be induced in the discrete entropy equation. The proposed FD methods furthermore are expected to work on general triangle/tetrahedral grids via a recent extension of the SBP framework to tensor-product spectral-element operators on collapsed coordinates [58]. In addition, while the current results consider single-block periodic domains, incorporating entropy-stable interface conditions for boundaries and inter-block coupling is possible via Simultaneous Approximation Terms [33, 59], as is typically done with SBP-compatible nonlinear flux methods.
Different avenues for future development remain. First, the schemes presented in this paper admit the possibility of using non-symmetric EC fluxes; such biased forms could be useful in the context of reducing oscillations around sharp gradient fields, and exploring this topic could be the subject of future research. Secondly, the methods’ extension to high-order stems naturally from employing high-order finite difference stencils to the differential definitions of the specialized averages; however, this will require extending the two-point algorithm for logarithmic means [47] to a multi-point format that is capable of effectively treating instances where the quotient formula becomes singular. In the meantime, however, Appendix A provides an alternate error-reducing treatment based on flux differencing for the narrow-width biased methods. Finally, the prospect of extending the current framework to general equations of state per the recent work of Aiello et al. [60] is to be explored.
Appendix A Reducing errors in the biased EC schemes via flux differencing
The current FD-EC methods would naturally extend to high-order upon employing high-order derivative formulas within the differencing as well as the FD definitions of the specialize averages. However, the latter requires the development of new multi-point algorithms for calculating the mean quantities while appropriately handling the risk of a singularity when evaluating the quotient expressions. Here, in the meantime, we propose a error-reducing extension of the narrow biased FD-EC schemes based on flux differencing.
$\frac{\mathcal{F}^{(n)}_{i+1/2}-\mathcal{F}^{(n)}_{i-1/2}}{h_{i}\cdot\Delta x}=\frac{1}{\Delta x}\left(\sum_{k\leq 0}2d_{i,i+k}^{\text{c}}\cdot\mathcal{F}^{(n)}\big{|}^{\alpha}_{\text{2pt}}(q_{i+k},q_{i})+\sum_{k>0}2d_{i,i+k}^{\text{c}}\cdot\mathcal{F}^{(n)}\big{|}^{\alpha}_{\text{2pt}}(q_{i},q_{i+k})\right)$ (94)
where
$\mathcal{F}^{(n)}\big{|}^{\alpha}_{\text{2pt}}(q_{a},q_{b})=\left[\begin{array}[]{c}\mathcal{F}^{(n)}_{\rho}\big{|}^{\alpha}_{\text{2pt}}\\ \\ \mathcal{F}^{(n)}_{\rho u_{m}}\big{|}^{\alpha}_{\text{2pt}}\\ \\ \mathcal{F}^{(n)}_{\rho E}\big{|}^{\alpha}_{\text{2pt}}\end{array}\right]=\left[\begin{array}[]{c}\overline{(\rho_{a},\rho_{b})}^{\log,n}\left((1-\alpha)\cdot u_{n,a}+\alpha\cdot u_{n,b}\right)\\ \\ \mathcal{F}^{(n)}_{\rho}\big{|}^{\alpha}_{\text{2pt}}\cdot\frac{(u_{m,b}+u_{m,a})}{2}+\delta_{mn}\cdot\left((1-\alpha)\cdot p_{b}+\alpha\cdot p_{a}\right)\\ \\ \mathcal{F}^{(n)}_{\rho}\big{|}^{\alpha}_{\text{2pt}}\cdot\left(\sum_{m=1}^{d}\frac{u_{m,a}u_{m,b}}{2}+\overline{(e_{a},e_{b})}^{H\log,n}\right)\\ +\ (1-\alpha)\cdot u_{n,a}p_{b}+\alpha\cdot u_{n,b}p_{a}\end{array}\right]\ .$ (95)
In the above, $d_{i,i+k}^{\text{c}}$ are the stencil coefficients associated with a central SBP operator evaluated at node $i$. While the fluxes are fully biased with respect to the advective velocity and the pressure, these quantities are counter-balanced and therefore the overall method and its fluxes can be interpreted as being “centered". As such, a central SBP operator in the prescription of the $d_{i,i+k}^{\text{c}}$ coefficients is still appropriate. The multi-point flux associated with the above is (see Eq. 3.9 in [41])
$\mathcal{F}^{(n)}\big{|}^{\alpha}_{\text{multi-pt}}=\sum_{k>0}\sum_{\ell\leq 0}2d_{i+\ell,i+k}^{\text{c}}\cdot\mathcal{F}^{(n)}\big{|}^{\alpha}_{\text{2pt}}(q_{i+\ell},q_{i+k})\ .$ (96)
Here we revisit the isentropic vortex and Sod shock tube test cases from Section 5 in order to survey the newly proposed methods; we focus on the backward biased ECBand symmetric ECSrenditions for brevity. The long-time entropy conservation of the new formulations are confirmed (see Figures 6(b) and 7(f)). Meanwhile, the convergence test on the isentropic vortex (see Figure 6(a)) shows that the current multi-point flux-differencing of the biased schemes are still limited to first order; however, the error constant is reduced from the nominal two-point second order method previously shown in Figure 1(a). Despite not being formally high-order, the current multi-point flux-differencing extension of the biased schemes is also seen to reduce oscillations in Sod shock tube results compared to the symmetric flux (see Figure 7). This property therefore motivates the potential utility of such schemes for flows exhibiting sharp gradients.
(a) Time evolution of entropy integral
(b) Time evolution of entropy integral
Figure 6: Isentropic vortex test for different discretizations using flux differencing: solid black lines are used for ECBand red for ECS. $\text{CFL}=0.01$.

### v2

@@ 4.1 On the kinetic-energy and pressure-equilibrium preservation of the new schemes
$\xrightarrow[]{\text{(2pt)}}\mathcal{F}^{(n)}_{\rho k,i+1/2}=\frac{1}{4}\left[(\tilde{\rho}u_{n}u_{m})_{i}(u_{m})_{i+1}+(\tilde{\rho}u_{n}u_{m})_{i+1}(u_{m})_{i}\right]$
We note that the convective term in the kinetic energy equation is in a quadratically split form which will result in local conservation of the quantities. Analogous expressions can be written for the biased stencils. Although kinetic energy preservation does not provide a rigorous nonlinear stability metric for compressible flows, incorporating its numerical consistency has been shown to greatly enhance solution quality—even for the class of entropy-conserving Tadmor fluxes [15, 32]. And as motivated in Eqs. (5) and (2.1) of Section 2, it forms a key consistency within the continuous entropy analysis.
Next, it is possible to also show that all the new FD-EC schemes have the Pressure Equilibrium Preservation (PEP) property [4, 31, 63] for the current context of a calorically perfect single fluid. In other words, they have the ability to preserve the equilibrium of velocity and pressure across an interface. The condition on the discretization of the momentum equation assuming constant velocity $U_{o}$ and pressure $P_{o}$ (see [63, 31]) is that the momentum flux is equal to
@@
4.2 On the damping effects of split-form stencil biasing
The schemes presented in the current section represent an array of different biasing and stencil dependencies, which are then expected to impact the resulting solution. Qualitatively understanding the impact of such choices is possible upon parameterizing the stencil biasing as a central component plus an artificial dissipation term,
$\delta^{(\omega)}\triangleq\delta^{c}+\omega\cdot\delta^{AD},\ \text{with}\ \omega\in[-1,1]\ \to\ \delta^{(\omega=0)}=\delta^{c},\ \delta^{(\omega=\pm 1)}=\delta^{\pm}\ .$ (94)
Note that the above is simply a rearrangement of Eq. (13), with $\omega$ controlling the level of biasing. Such a re-writing, however, allows one to highlight the anticipated damping effects inherent within the biased split-forms, as will be shown next.
The overall group of proposed finite-difference compatible entropy-conserving (FD-EC) methods can be expressed in semi-discrete residual form as
$d_{t}q=\alpha\cdot\mathcal{R}^{(\omega)}+(1-\alpha)\cdot\mathcal{R}^{(-\omega)},\ \text{where}\ \mathcal{R}^{(\omega)}=-(\mathcal{C}^{(\omega)}+\mathcal{P}^{(\omega)})\ ,$ (95)
with the convective and pressure terms summarized in Table 3 and depending on the biasing parameter $\omega$. Equation (95) is thus a two-parameter family of schemes based on the pair $\alpha,\omega]$. The centered wide stencils are recovered for the parameter pair $(\alpha,\omega=0)$; the centered narrow stencils are recovered for $(\alpha=1/2,\omega\neq 0)$; and, the “fully" biased stencils (i.e., at least relative to velocity and pressure) are recovered for $(\alpha=\pm 1,\omega=\pm 1)$.
$\bm{\mathcal{C}}_{\rho}^{(n,\omega)}$ $\delta^{(\omega)}_{x_{n}}(\tilde{\rho}^{(-\omega)}u_{n})$
$\bm{\mathcal{C}}_{\rho u_{m}}^{(n,\omega)}$ $\frac{1}{2}\left[\delta_{x_{n}}^{(\omega)}(\tilde{\rho}^{(-\omega)}u_{n}u_{m})+\tilde{\rho}^{(-\omega)}u_{n}\delta_{x_{n}}^{(-\omega)}u_{m}+u_{m}\delta_{x_{n}}^{(\omega)}(\tilde{\rho}^{(-\omega)}u_{n})\right]$
$\bm{\mathcal{C}}_{\rho e}^{(n,\omega)}$ $\delta^{(\omega)}_{x_{m}}(\tilde{\rho}^{(-\omega)}u_{n}\tilde{e}^{(-\omega)})$
$\bm{\mathcal{P}}_{\rho u_{m}}^{(n,\omega)}$ $\delta^{(-\omega)}_{x_{m}}p$ if $n=m$, otherwise 0
$\bm{\mathcal{P}}_{\rho e}^{(n,\omega)}$ $p\delta^{(\omega)}_{x_{n}}u_{n}$
Table 3: A kinetic-energy and pressure-equilibrium preserving finite-difference discretization of Euler derivatives assuming SBP-compatible stencils. The auxiliary variables are $\tilde{\rho}^{(\omega)}\triangleq\frac{\delta_{x_{n}}^{(\omega)}\rho}{\delta_{x_{n}}^{(\omega)}\log\rho}\equiv\overline{\rho}^{(\log,n,\omega)}$ and $\tilde{e}^{(\omega)}\triangleq\frac{\delta^{(\omega)}\log e^{-1}}{\delta^{(\omega)}e^{-1}}\equiv\overline{e}^{(H\log,n,\omega)}$. Centered wide stencil method is recovered for $\omega=0$.
Leveraging Eq. (94), the split-form biasing associated with the discretized residual can be decomposed into “central" and “dissipation" components for a given bias parameter $\omega$, $\mathcal{R}^{(\omega)}=(\mathcal{R}^{(\omega),c}+\mathcal{R}^{(\omega),AD})$:
$-\mathcal{R}_{n}^{(\omega),c}$ $=\left[\begin{array}[]{l}\delta_{x_{n}}^{c}\tilde{\rho}^{(-\omega)}u_{n}\\ \\ \frac{1}{2}\left[\delta_{x_{n}}^{c}\tilde{\rho}^{(-\omega)}u_{n}u_{m}+\tilde{\rho}^{(-\omega)}u_{n}\delta_{x_{n}}^{c}u_{m}+u_{m}\delta_{x_{n}}^{c}\tilde{\rho}^{(-\omega)}u_{n}\right]+\delta_{x_{n}}^{c}p\\ \\ \delta_{x_{n}}^{c}\tilde{\rho}^{(-\omega)}u_{n}\tilde{e}^{(-\omega)}+\frac{1}{2}\left[u_{m}\delta_{x_{n}}^{c}\tilde{\rho}^{(-\omega)}u_{n}u_{m}+\tilde{\rho}^{(-\omega)}u_{n}u_{m}\delta_{x_{n}}^{c}u_{m}\right]+(u_{m}\delta_{x_{n}}^{c}p+p\delta_{x_{n}}^{c}u_{m})\end{array}\right]$ (101)
$-\mathcal{R}_{n}^{(\omega),AD}$ $=\left[\begin{array}[]{l}\omega\cdot\delta_{x_{n}}^{AD}\tilde{\rho}^{(-\omega)}u_{n}\\ \\ \omega\cdot\left[\frac{1}{2}\delta_{x_{n}}^{AD}\tilde{\rho}^{(-\omega)}u_{n}u_{m}+\frac{1}{2}\left(u_{m}\delta_{x_{n}}^{AD}\tilde{\rho}^{(-\omega)}u_{n}-\tilde{\rho}^{(-\omega)}u_{n}\delta_{x_{n}}^{AD}u_{m}\right)-\delta_{x_{n}}^{AD}p\right]\\ \\ \omega\cdot\left[\delta_{x_{n}}^{AD}\tilde{\rho}^{(-\omega)}u_{n}\tilde{e}^{(-\omega)}+(\tilde{\rho}^{(-\omega)}u_{n}u_{m}\delta_{x_{n}}^{AD}u_{m}-u_{m}\delta_{x_{n}}^{AD}\tilde{\rho}^{(-\omega)}u_{n}u_{m})+(p\delta_{x_{n}}^{AD}u_{m}-u_{m}\delta_{x_{n}}^{AD}p)\right]\ .\end{array}\right]$ (107)
From the above, one can anticipate the narrow split forms to produce an additional damping effect in the density and pressure fields compared to the wide-stencil configuration of $\omega=0$. This will occur even though the overall formulation is kinetic-energy-preserving and entropy-conserving, and it may be more noticeable in flows featuring sharp gradients. The narrow symmetric implementation (i.e., $(\alpha=1/2,\omega\neq 0)$ will lack additional pressure damping compared to the biased renditions. Namely, the narrow symmetric scheme lacks $\delta_{x_{n}}^{AD}p$ in the momentum equation as well as $(p\delta_{x_{n}}^{AD}u_{m}-u_{m}\delta_{x_{n}}^{AD}p)$ in the total energy equation. Such nuances highlight the distinctness of the respective schemes, which is further explored in the following numerical evaluations.
@@ 5 Numerical results
The new EC schemes are here tested with a variety of cases (i.e., the 2D isentropic vortex, the 1D density wave, the 1D Sod shock tube, and the 3D Taylor–Green vortex) in order to verify the validity of the theoretical results. In particular, we consider the new scheme based on the wide stencil presented in Eq. (77), here referred to as ECW, and three schemes from the family identified in Eq. (84), that with $\omega=0$, $\omega=1$, and $\omega=0.5$ are labeled ECB, ECF, and ECS, respectively. Notably, ECSis analogous to the scheme of Ranocha in the scenario of a two-point second order stencil. A summary of the schemes and their respective numerical fluxes is presented in Table 4.
@@
Despite having demonstrated the capability to obtain high-order formulations of the presented EC schemes, the methods will all be tested with their low-order version, employing first and second order accurate differential operators. This is due to the singularity problem that occurs when evaluating the logarithmic mean for almost identical values. This challenge is handled in the case of two-point means by employing suitable expansion approximations such as the one famously introduced by Ismail and Roe [58]. Other similar approaches have been recently employed to produce schemes that are asymptotically entropy conserving [33, 34, 35]. However, a generalization of the procedure for multiple-point filters has yet to be developed.
The results presented in this section have been generated based on the flux forms of the FD schemes in the code STREAmS-2 [65]. Density, momentum and total energy are employed as the primary variables. All time integrations have been performed using the classical explicit fourth-order Runge-Kutta (RK4) scheme. In order to analyze the conservation properties of the schemes, we introduce the notation of using the angular brackets to indicate integration over the spatial domain $\langle\phi\rangle=\int_{\Omega}\phi\,\mathrm{d}\Omega$ and using $\epsilon$ for the nondimensionalization $\epsilon_{\phi}=\left(\langle{\phi}\rangle-\langle{\phi_{0}}\rangle\right)\left/\left|\langle{\phi_{0}}\rangle\right|\right$., in which $\phi_{0}$ indicates the values at the starting time. For all the simulations, the value for the normalized specific heat capacity at constant volume is $c_{v}=1$ and $\gamma=1.4$.
Ref. $\mathcal{F}_{\rho}^{(n)}$ $\mathcal{F}_{\rho u_{m},i+1/2}^{(n)}$ $\mathcal{F}_{\rho E}^{(n)}$ ECB new $\overline{(\rho_{i},\rho_{i+1})}^{(\log)}\,u_{n,i}$ $\mathcal{F}_{\rho,i+1/2}\overline{u_{m}}\,+p_{i+1}$ $\mathcal{F}_{\rho}\overline{(e_{i},e_{i+1})}^{(H\log)}\,+\mathcal{F}_{\rho}\,\sum_{m}\frac{u_{m,i}u_{m,i+1}}{2}+p_{i+1}u_{n,i}$ ECF new $\overline{(\rho_{i},\rho_{i+1})}^{(\log)}\,u_{n,i+1}$ $\mathcal{F}_{\rho}\overline{u_{m}}\,+p_{i}$ $\mathcal{F}_{\rho}\overline{(e_{i},e_{i+1})}^{(H\log)}\,+\mathcal{F}_{\rho}\,\sum_{m}\frac{u_{m,i}u_{m,i+1}}{2}+p_{i}u_{n,i+1}$ ECS [15] $\overline{(\rho_{i},\rho_{i+1})}^{(\log)}\,\overline{u_{n}}\,$ $\mathcal{F}_{\rho}\overline{u_{m}}\,+\overline{p}\,$ $\mathcal{F}_{\rho}\overline{(e_{i},e_{i+1})}^{(H\log)}\,+\mathcal{F}_{\rho}\,\sum_{m}\frac{u_{m,i}u_{m,i+1}}{2}+\overline{\overline{(p,u)}}$ ECW new $0.5\overline{(\rho_{i-1},\rho_{i+1})}^{(\log)}\,u_{n,i}$ $\mathcal{F}_{\rho}\overline{u_{m}}\,+\overline{p}\,$ $0.5\overline{(\rho_{i-1},\rho_{i+1})}^{(\log)}\,u_{n,i}\overline{(e_{i-1},e_{i+1})}^{(H\log)}\,$ $0.5\overline{(\rho_{i},\rho_{i+2})}^{(\log)}\,u_{n,i+1} +$$0.5\overline{(\rho_{i},\rho_{i+2})}^{(\log)}\,u_{n,i+1}\overline{(e_{i},e_{i+2})}^{(H\log)}\, +$$\mathcal{F}_{\rho}\,\sum_{m}\frac{u_{m,i}u_{m,i+1}}{2}+\overline{\overline{(p,u)}}$
@@
Table 4: Numerical fluxes for the schemes used for the tests. The fluxes are evaluated along direction $n$ at location $(i+1/2)$. All interpolation operations are implicitly interpreted to be in the same direction $n$.
@@
The isentropic vortex flow is an exact solution of the inviscid compressible flow equations and as such it can be used to evaluate the accuracy of the numerical methods. In Fig. 1(a) the error on momentum along the $x$ axis is displayed as a function of $N$ number of nodes in each direction. The error is evaluated as the maximum norm of the difference between the computed and the exact solution for momentum along the $x$ direction $\|\rho u-(\rho u)_{ex}\|_{\infty}$, evaluated at a time $t=0.01$ with $\text{CFL}=0.01$. The analysis shows a difference between the first-order accurate (ECFand ECB) and the second-order accurate (ECSand ECW) methods. The reason behind this behavior is easily explained by the fact that ECFand ECBuse an asymmetrical numerical derivative for the discretization of the pressure term and an asymmetrical interpolator for velocity as well; on the other hand both ECSand ECWresult in central schemes. Identical trends are witnessed upon inspecting the error of other quantities (not shown), such as the primitive variables $(\rho,u,T)$.
@@
Figure 2: Density wave test for different discretizations: black lines are used for ECB, red for ECS, green for ECF, blue ECW; dashed magenta line shows the exact solution for density. $\text{CFL}=0.01$, final solutions at $T=40$.
It can be shown theoretically (see Section 4.1) that all of the current schemes have the PEP property. The simulation of a travelling density wave is well suited to numerically test the PEP property of the schemes. The initial conditions for the test are
@@
with the domain [L,L][-L,L] with $L=1$ discretized in $61$ points and periodic boundary conditions. The Courant number has been chosen to be $\text{CFL}=0.001$ to prevent errors on entropy conservation due to the temporal integrator, and the simulation has been carried out until a final time $T=40$, which corresponds to approximately 20 cycles of the wave through the domain.
@@
The simulation is carried out until $T=0.1$ with $\text{CFL}=0.01$ based on velocity $u_{\infty}=\sqrt{p_{U}/\rho_{U}}$. For the discretization of the interval [0.25,0.25][-0.25,0.25], $101$ nodes are employed, however a much larger domain has been used to prevent the periodic boundary conditions to have any effect on the solution.
In order to appreciate the difference between the various schemes, Fig. 3 shows the evolution in time of the $L_{2}$ norm of the difference between the computed solution and the analytical one in the interval [0.25,0.25][-0.25,0.25]. The analytical solutions has been obtained with an exact Riemann solver. While the ECSscheme seems to attain better results considering velocity (Fig. 3(b)), this is no longer the case when other variables are examined. In fact, at least one of the biased schemes, ECBand ECF, appears to be outperforming the symmetric schemes; this is especially evident in the case of density (Fig. 3(a)).
To evaluate the spatial distribution of the error, in Fig. 4, the solutions at time $T=0.1$ obtained with the various EC schemes are compared with the analytical one. As expected, all the schemes are entropy conserving (Fig. 4(f)) and generate spurious oscillations in the solution. As visible in Fig. 4(c), none of the schemes are able to reproduce the correct behavior for entropy in the region leading up to shock $x\in[0.1,0.2]$, and ECWseems to present more oscillations. This behavior has been observed before for the scheme of Ranocha [31] (here ECS) in [41]. This region is also one in which the oscillations are stronger for all the schemes when other variables are considered (e.g., density, velocity, temperature, pressure).
On the other hand, in the region $x\in[0,0.1]$ the asymmetric schemes ECBand ECFpresent oscillations with smaller amplitude when compared to the central schemes ECWand ECS. This is particularly evident when inspecting velocity in Fig. 4(b), for which a larger spike is present near $x=0$ for the ECSscheme. Such behavior is supported by the presence of additional damping terms in the density and pressure fields as suggested by Eqs. (101) and (107). Despite being kinetic-energy-preserving and entropy-conserving, the split-form biasing inherently provides a low level of regularization compared to the centralized configurations. This has also been observed in previous work for biased split-forms [49] and further motivates interest in studying these schemes for use in flows exhibiting sharp gradients.
While the focus of this work is on EC schemes, Appendix C demonstrates how they can be the base to Entropy Stable (ES) schemes through the addition of a specialized dissipative term. These ES schemes are tested again on Sod’s test, to evaluate their ability to suppress oscillations in the presence of strong gradients.
@@
(d) Pressure
Figure 3: Sod test for different discretizations: black lines are used for ECB, red for ECS, green for ECF, blue ECW. $\text{CFL}=0.01$. Evolution in time of the $L_{2}$ norm of the difference between computed solution and analytical one.